OpenStudy (anonymous):

A box has dimensions of 17 inches long, 1.3 feet wide, and 8 inches high. What is the volume of the box? The formula for the volume is V = l • w • h. 2121.6 cubic inches (in3) 176.8 cubic inches (in3) 14.7 cubic inches (in3) 1.2 cubic inches (in3)

OpenStudy (anonymous):

Well if V=l*w*h, just plug in the numbers. V=17*1.3*8. Put that sucker in a calculator and you get 176.8 inches cubed :)

OpenStudy (campbell_st):

well before you calculate you need all measurements in the same units 1.3 ft = ? inches...?

OpenStudy (anonymous):

Oh! I didn't even notice that!

OpenStudy (anonymous):

15.6 inches

OpenStudy (campbell_st):

well that now is fine... so multiply the 3 measurements to get the volume in cubic inches

OpenStudy (anonymous):

so its going to be 2121.6

OpenStudy (anonymous):

i have more questions

OpenStudy (campbell_st):

well just post them in the normal way and there are lots of people who can help